OpenStudy (anonymous):

Mrs. julie's classroom the ratio of boys to girls 12 to 8 if the elementary school has a population of 400 students how many would you expect to be girls. A. 3200 b. 600 c. 260 d. 160

OpenStudy (anonymous):

12/8 = 3/2......240/160 = 3/2...girls = 160